OpenStudy (anonymous):

does 2/(n+1) converge/diverge?

OpenStudy (amistre64):

neither id say ....

OpenStudy (amistre64):

2/(n+1) is just an expression that needs to be applied to something in order to have some sort of convergence or divergence ...

OpenStudy (amistre64):

the sequence the is defined by this has a limit that goes to zero; since the bottom goes large the top goes small

OpenStudy (amistre64):

not the top perse; but the value :)

OpenStudy (anonymous):

welll there is the summation, I didn't put it in. \[\sum_{n = 0}^{\infty}\]2/(n+1)

OpenStudy (amistre64):

the partial sums .... i cant tell at the moment tho

OpenStudy (amistre64):

is that for the sequence of the partial sums?

OpenStudy (anonymous):

which test should I use?

OpenStudy (anonymous):

no, it is not for partial sums.

OpenStudy (amistre64):

then id just play it out as is; cant recall any details for the "tests" at the moment. the limit of 2/(n+1) as n approaches infinity is 0 since the bottom outweighs the top for large values of n

OpenStudy (amistre64):

the sequence is really a set of the range of the function. So whatever the function behaves for large values of n; it either goes to zero and converges, or it doesnt and diverges ...

OpenStudy (anonymous):

ok, thank you

OpenStudy (amistre64):

youre welcome :)

OpenStudy (phi):

Your summation is \[\sum_{n=0}^{\infty}\frac{2}{n+1}= 2\sum_{n=1}^{\infty}\frac{1}{n}\] which is 2 times the harmonic series. This series is famous for not converging. You can show this using the integral test.